Habanero  Jigsaw  Infospheres  JavaSoft  Netscape
Module  collaboratized  Resource  dapplet→  Java Bean  LiveWire app.
applet  DJINN  server
Port/Channel Java socket  any HTTP  portlet→  RMI  custom?
carrier  mailbox
Message  Marshalled Event Pickled  any object  Serialized Object JavaScript
or Action  Resource  bytesteam
Compute-Web Star topology  2-node  any topology
Runtime  Collaboratory  Java HTTP  dapplet/DJINN Jeeves (Java  community or
server  server  manager?  server)  enterprise system
User Interface AWT  Forms  visual authoring? HotJava  Navigator
Coordination instantaneous  client-server  asynchronous  CORBA  multi-server
broadcast  multi-server
Persistency Resource Store flat file? JDBC LiveWire→ DB
Publication  javadoc